Circular No. 6058

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994Y IN NGC 5371 William Wren, McDonald Observatory, reports his visual discovery of

an apparent supernova in NGC 5371 (α = 13^{h}55^{m}.7, δ = +40^{o}28′, equinox 2000.0), located ∼ 28′′ west and 14′′ north of the galaxy’s nucleus. Wren provides the following magnitude estimates as seen through a 0.9-m telescope: July 31.18 UT, [16.5; Aug. 19.15, 15.0.

Y. Paik, A. V. Filippenko, R. R. Treffers, and S. D. Van Dyk, University of California at Berkeley; and M. W. Richmond, Princeton University, report that R-band images of NGC 5371 were obtained in August as part of the Leuschner Observatory Supernova Search. On Aug. 12, SN 1994Y is 2.7 mag fainter than the comparison star 61′′ west and 5′′ south of it, barely below the threshold for automatic detection. On Aug. 16, it is perhaps 2 mag fainter than the comparison star, but the image quality is very poor. Images obtained on Aug. 3, 6, and 9 show little or nothing due to clouds. SN 1994Y is measured as 34′′ west and 11′′ north of the galaxy’s nucleus.

SUPERNOVA 1993J IN NGC 3031 R. W. Hanuschik, University of Tübingen; and M. Werger, University

of Bochum, report: “We have performed optical spectroscopy (dispersion 20.0 nm/mm) of SN 1993J on June 6–8 and Aug. 1–3 with the 2.2-m telescope (+ CAFOS Faint Object Spectrograph) of the German-Spanish Observatory at Calar Alto. We confirm the outburst of broad Hα emission as reported on IAUC 6005; if due to pure Hα emission, its total base width is ∼ 20000 km/s. We do not exclude, however, the alternative possibility of blended Hα and He I 667.8-nm emission, in which case the half base width of each line is ∼ 5000 km/s. The intensity of the emission feature is approximately the same as that of [O I] at 630.0 and 636.3 nm in June, and higher than that in August; it is also substantially higher than [Ca II] at 729.1 and 732.3 nm for June and August. If this feature is a blend of Hα and He I, then both lines show about the same intensity for each date. The emission feature is likely to be the result of the interaction of the expanding ejecta with a dense shell in the progenitor’s circumstellar wind.”

SUPERNOVA 1994W IN NGC 4041 Visual magnitude estimates by L. Szentaskó, Veresegyház, Hungary:

Aug. 8.83 UT, 13.5; 10.84, 14.0; 12.83, 13.9; 16.83, 14.0; 18.83, 13.8.

1994 August 19 Daniel W. E. Green